OpenStudy (anonymous):

Enrollment in a business class decreased from 25 to 15 students. What was the percent of decrease in enrollment?

OpenStudy (anonymous):

40%?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \% decrease = {\Large (}{Decrease \over Original} \cdot 100{\Large )}\%}\)

OpenStudy (anonymous):

o.O using that i got 60 percent... but that cant be right

OpenStudy (anonymous):

if it dropped 60% it would be less than half of 25 and 15 is defiantly greater than half of 25

OpenStudy (asnaseer):

ryo: your original answer of 40% is correct.

OpenStudy (anonymous):

thanks i used my fingers XD

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

decrease = 25 - 15 = 10 so percentage decrease = (10/25) * 100 % = 40%

OpenStudy (anonymous):

ooooh ok thanks

OpenStudy (asnaseer):

yw :)